The near point of an eye is 111 cm. A corrective lens is to be used to allow this eye to focus clearly on objects 25.0 cm in front of it. What should be the focal length of this lens?
What is the power of the needed corrective lens (in diopters)? Do not enter units.
Answer #1

Here ,

focal length is f

object distance , do = -25 cm

image distance ,di = -111 cm

Using lens formula

1/f = 1/v - 1/u

1/(f) = 1/(-111) + 1/25

f = 32.3 cm

the focal length of the lens is 32.3 cm

power of the corrective lens = 100/f

power of the corrective lens = 100/32.3

power of the corrective lens = 3.1 D

the power of the corrective lens is 3.1 D